# Dynamic Reconstruction of Hand-Object Interaction with Distributed Force-aware Contact Representation

### v1

@@ Abstract
We present ViTaM-D, a novel visual-tactile framework for dynamic hand-object interaction reconstruction, integrating distributed tactile sensing for more accurate contact modeling. While existing methods focus primarily on visual inputs, they struggle with capturing detailed contact interactions such as object deformation. Our approach leverages distributed tactile sensors to address this limitation by introducing DF-Field. This distributed force-aware contact representation models both kinetic and potential energy in hand-object interaction. ViTaM-D first reconstructs hand-object interactions using a visual-only network, VDT-Net, and then refines contact details through a force-aware optimization (FO) process, enhancing object deformation modeling. To benchmark our approach, we introduce the HOT dataset, which features 600 sequences of hand-object interactions, including deformable objects, built in a high-precision simulation environment. Extensive experiments on both the DexYCB and HOT datasets demonstrate significant improvements in accuracy over previous state-of-the-art methods such as gSDF and HOTrack. Our results highlight the superior performance of ViTaM-D in both rigid and deformable object reconstruction, as well as the effectiveness of DF-Field in refining hand poses. This work offers a comprehensive solution to dynamic hand-object interaction reconstruction by seamlessly integrating visual and tactile data. Codes, models, and datasets will be available.
footnotetext: * indicates equal contributions.
§ Cewu Lu is the corresponding author.
@@ 1 Introduction
Human manipulates objects with tactile feedback. Reconstructing the manipulation process of hand-object interaction is an important task that can benefit many downstream tasks, such as VR/AR, robot imitation learning, and human behavior understanding. Previous works [31, 34, 4, 27, 12] on dynamic hand-object interaction reconstruction are primarily visual-only. Visual-only methods can recover the global geometry and poses of hand and object, but they struggle with contact details such as object deformation due to a lack of information near contact areas.
Recently, visual-tactile methods [29, 23, 28] have increasingly drawn attention in hand-object reconstruction tasks, with the development of tactile sensing techniques [24, 16, 37, 21], which can supplement the perceptual ability near contact areas (Fig. 1.a). Among the fast-developing tactile sensors, distributed tactile sensors (Fig. 1.b) [16, 36, 24], which are wearable or attachable to a hand and allow collecting tactile information in a natural human-like hand-object interaction process, are more promising in human manipulation data collection than other types of tactile sensors [25, 37, 15, 21]. However, how to integrate such sensors with visual perception to reconstruct the states of hand-object interaction is seldom explored.
@@
Figure 1: (a): The relationship between tactile information and contact geometry. Grasping different bowls with the same hand poses shows that the distributed tactile arrays can capture occlusion contacts and object states. (b): Different types of distributed tactile sensors proposed in previous works (1). [16], (2). [36], (3). [24].
@@
Due to the conformality between the distributed tactile sensor and the hand, the hand geometry will not be significantly influenced during manipulation. Therefore, we can use the same visual-only approach to reconstruct the hand-object state and integrate the tactile information to refine the result. Based on this observation, we propose a novel Distributed Force-aware contact representation, DF-Field, and a Visual-Tactile Manipulation reconstruction framework with Distributed tactile sensing, ViTaM-D. DF-Field models the contact by considering both kinetic and potential energy in hand-object interaction, which allows modeling object deformation. With this representation, ViTaM-D first reconstructs the hand-object interaction with visual-only observations by the proposed Visual Dynamic Tracking network, VDT-Net, and refines the contact details with DF-Field via a Force-aware hand-pose Optimization process, FO. Such design allows us to inherit the wisdom of fast-growing visual-only hand-object reconstruction methods [34, 4, 27, 12] and easily integrate the tactile information into an existing motion capture or estimation system.
@@
To train the VDT-Net, we need large-scale hand-object interaction datasets since previous datasets [3, 33, 9] of hand-object interactions mostly cover rigid or articulated object manipulation, failing to contain deformable objects. And they generally do not provide accurate tactile readings. Therefore, aside from the public dataset, DexYCB [3], which we adopt to benchmark on rigid objects with common baselines, we also create a new dataset, HOT dataset, to fully benchmark our method on deformable object reconstruction. The dataset is built with ZeMa [8], a high-precision physics-based simulation environment that supports penetration-free frictional contact modeling with finite element method (FEM) to model the object deformation. The HOT dataset contains 600 sequences of hand-object interaction, with 30 deformable objects from 5 different categories and 8 camera views for each sequence.
@@
To evaluate the method, we compare our proposed ViTaM-D with previous state-of-the-art methods gSDF [6] and HOTrack [4] on DexYCB. Extensive experiments proved that our method realizes great improvements in the previous works both quantitatively and qualitatively. Besides, we also prove the great capability of our method in tracking deformable objects on the HOT Dataset and the effectiveness of the FO optimization in refining hand poses from penetrations and bad contact states.
@@
(1) A visual-tactile learning framework, ViTaM-D. It contains a visual-only dynamic tracking network, VDT-Net, for reconstructing hand-object interactions and a force-aware optimization process, FO, to integrate the tactile information into reconstruction refinement based on a novel distributed force-aware contact representation, DF-Field.
@@ 2 Related Work
These works focus more on visual-only inputs. However, due to occlusion between the hand and object during the interaction, visual perception usually lacks information near the contact areas, and such information cannot necessarily be mitigated by cross-frame feature fusion. Therefore, tactile perception comes to supplement the near-contact information. Zhang et al. [38] utilized a tactile glove [24] to track object movements. However, it focuses more on the dynamic object trajectory rather than contact geometry. Works such as [23, 28] employed optical tactile sensors, trained the model by synthetic data, and applied to rigid object geometry reconstruction. Later, VTacO [29] extended this line of research by using optical tactile sensors to capture object geometry, including deformation, providing a more comprehensive representation of hand-object interactions. Unlike these works, the proposed ViTaM-D inherits the merits from both worlds; it leverages the advanced techniques of visual-only reconstruction and incorporates the distributed tactile readings to enhance the local contact details.
@@ 3 Force-aware Contact Representation
During manipulation, forces from the hand change both hand and object states. To fully capture dynamic contact behaviors, the representation must encode both contact locations and forces. We introduce a distributed force-aware contact representation, DF-Field (Sec. 3.1), and apply it using distributed tactile sensors (Sec. 3.2).
@@ 3.1 DF-Field Representation
Without loss of generality, we take the object-centric perspective to describe a hand-object interaction process, where the object is assumed to be fixed in the origin point, and the hand moves around the object. In this way, by ignoring the gravitational potential energy, the contact dynamics in a manipulation process are driven by the Relative Potential Energy, which is the resulting energy of the kinetic energy and deformation potential energy of the hand and object. Besides, we assume a virtual Barrier Energy lies between the hand and object contact surface to prevent penetration. With an established point pair $i,j$ of object vertex $\mathcal{V}_{i}^{o}$ and hand vertex $\mathcal{V}_{j}^{h}$, and the Euclidean distance of the point pair $l_{ij}=\left\lVert\mathcal{V}_{i}^{o}-\mathcal{V}_{j}^{h}\right\rVert_{2}$, we define the two energy terms as follows.
@@
This barrier energy aims to push away the point pair when $\kappa$ is low, thus avoiding penetration issues between the hand and object. The function is defined in this way not only to ensure repulsion becomes smaller when the distance is larger but also to remain smooth for optimization.
@@
To note, though $\kappa$ is strongly correlated to the tactile readings, DF-Field can also be tactile-independent, as long as we empirically set the exerted force instead of physics-based. In this way, DF-Field can work with visual-only methods. Different setups of force are discussed in Sec. 6.5.
@@ 3.2 DF-Field with Distributed Tactile Sensors
For optimization, we connect only the hand keypoints to object vertices, reducing computational load while enhancing regional interaction information. The force in each region is calculated by averaging the tactile readings $\mathcal{M}^{j}$, and by the definition of the energy, we can obtain that $\kappa$ is the difference of the exerted force over the distance $l_{ij}$. In practice, we approximate it by dividing the force by the distance:
@@
We will discuss the difference between choosing keypoints or all-hand vertices for optimization in Sec. Appendix B.2.
@@ 4.1 Overview
We address the problem of 4D tracking and dynamic reconstruction of hand-object interactions from a visual-tactile perspective and propose a novel method named ViTaM-D. We assume that the visual input consists of a live stream of 3D point clouds, denoted as $\{\mathcal{P}_{t}\in\mathbb{R}^{N\times 3}\}_{t=1}^{n}$, representing hand-object interactions derived from single-view depth images, where $N$ is the number of input point. The tactile input $\{\mathcal{M}_{t}\}_{t=1}^{n}$ contains distributed tactile sensor readings for each frame.
ViTaM-D tracks hand-object states in two stages: Visual Dynamic Tracking (VDT-Net) and Force-aware Optimization (FO). VDT-Net uses visual-only inputs to establish flow features and dynamically reconstruct the hand and object, capturing global hand-object geometry but struggling with sparse contact region details. To address this, the second stage, FO, incorporates tactile information to improve contact accuracy and reduce penetration issues. The overall pipeline is given in Fig. 2.
Figure 2: ViTaM-D pipeline. (a): The overview of our pipeline, including the flow prediction module for visual feature extraction and flow estimation, the Hand and Object Decoders, and Force-aware hand-pose Optimization, FO. (b): The Object Decoder to reconstruct the object mesh. (c): The Hand Decoder for estimating hand parameters and reconstructing based on the MANO model.
4.2 Visual Dynamic Tracking of hand and object, VDT-Net
In the first stage, the VDT-Net first predicts the flow from the last to the current frame with a flow prediction module and extracts a fused visual feature from the current frame’s flow and point cloud. This feature will then be forwarded into Object Decoder and Hand Decoder to reconstruct the object and hand geometry, respectively. We use the Signed Distance Field (SDF) to model the objects and the MANO model [22] for the hand.
@@ 4.2 Visual Dynamic Tracking of hand and object, VDT-Net
In $\mathcal{F}_{f}$, we first perform a Cartesian product of the two extracted features, yielding a tensor of size $N\times N\times 2d$. This tensor is fed into a 3-layer MLP to obtain $\mathcal{C}_{v}$, which is then used in two ways. First, it passes through a 2D convolutional layer for downsampling to obtain $p_{c}\in\mathbb{R}^{N\times N}$, representing the matching probability of each point between two frames. Second, $\mathcal{C}_{v}$ is sent through a softmax function and a one-layer MLP to downsample, resulting in $\mathcal{C}_{c}\in\mathbb{R}^{N}$, indicating whether the points in the first frame are matched in the second frame. Thus, the final matching probability matrix $p_{m}\in\mathbb{R}^{N\times N}$, which describes the correspondence likelihood between the two point sets, can be computed as:
$p_{m}=p_{c}\times\mathcal{C}_{c}$ (6)
After estimating the matching probability, we compute the disparity of two point cloud sets $\mathcal{D}\in\mathbb{R}^{N\times N\times 3}$, with $\mathcal{D}_{ij}=\mathcal{P}_{t}^{i}-\mathcal{P}_{t-1}^{j}$, and concatenate the disparity with $p_{m}$. The concatenated tensor is then fed into four 2D convolutional layers with batch normalization and a softmax function to obtain the disparity feature $F_{d}\in\mathbb{R}^{N\times d^{\prime}}$. Finally, we use PointNet++ [20] to regress the flow $f_{t-1\rightarrow t}\in\mathbb{R}^{N\times 3}$. The flow prediction results will be discussed in Appendix B.1.
@@
Additionally, we use another PointNet++ $I^{\prime}(\cdot)$ to extract the visual correspondence feature $F_{t}^{f}$ from $\mathcal{P}_{t-1}$ with $f_{t-1\rightarrow t}$ added. $F_{t}^{f}$ and $F_{t}$ are then forwarded to a transformer fusion module to obtain the final visual feature $F_{v}$, corresponding to the current frame’s point cloud $\mathcal{P}_{t}$. Specifically, the transformer fusion module first uses a self-attention module to encode the point cloud features from both frames. After applying positional embedding between the feature and its point cloud, a cross-attention module fuses the two features and outputs the final visual feature. This fusion strategy considers both the 3D static information of the current frame and the corresponding feature extracted from the flow.
@@
Object Decoder. Based on the final visual feature $F_{v}$, the Object Decoder follows two steps: Feature Scattering & Sampling, and SDF Decoding, to obtain the SDF predictions and the object mesh using Marching Cubes algorithm [17]. We follow the design in ConvOccNet [19], which scatters the feature into volume with the resolution $D$, feeds it into a 3D-UNet, and finally uses a 5-layer MLP to predict the signed distance for every point.
Hand Decoder. For hand tracking and reconstruction, we adopt the parametric MANO hand model [22], which uses $\beta\in\mathbb{R}^{10}$ for hand shape, and $\theta\in\mathbb{R}^{51}$ for hand poses.
Based on the extracted final visual feature $F_{v}$, and the point cloud $\mathcal{P}_{t-1}$ of the last frame as input, we inherit a voting mechanism in PVN3D [14] to predict the hand joint positions $\mathcal{J}_{h}\in\mathbb{R}^{21\times 3}$ of the current frame. Besides predicting the translation offset $\mathcal{O}_{t}\in\mathbb{R}^{21\times 3}$, we also predict the initial position of the keypoint by estimating a probability matrix $T_{t-1}\in\mathbb{R}^{N\times 21}$ and multiply it with the former frame point cloud $\mathcal{P}_{t-1}$. The final prediction of the hand joint locations $\mathcal{J}_{h}$ is:
$\mathcal{J}_{h}=\mathcal{O}_{t}+\mathcal{P}_{t-1}\times T_{t-1}$ (7)
After obtaining hand joint locations $\mathcal{J}_{h}$, we use the inverse kinematics method [39] to find the joint poses $\theta$ with a template hand shape $\beta$, and we can easily reconstruct the hand mesh through a differentiable MANO layer.
Loss Function. We train the network in an end-to-end way, and optimize the framework parameters with three loss terms:
$\mathcal{L}=\lambda_{f}\mathcal{L}_{flow}+\lambda_{S}\mathcal{L}_{SDF}+\lambda_{H}\mathcal{L}_{Hand},$ (8)
@@
where flow loss is defined by two Chamfer distance terms $\mathcal{L}_{flow}=CD(\mathcal{P}_{t-1}^{{}^{\prime}},\mathcal{P}_{t})+CD(\mathcal{P}_{t}^{{}^{\prime\prime}},\mathcal{P}_{t-1})$, $\mathcal{P}_{t-1}^{{}^{\prime}}=\mathcal{P}_{t-1}+f_{t-1\rightarrow t}$ is the forward-shifted point cloud and $\mathcal{P}_{t}^{{}^{\prime\prime}}=\mathcal{P}_{t}+f_{t\rightarrow t-1}$ is the backward-shifted point cloud. SDF loss is $\mathcal{L}_{SDF}=|s-s^{*}|$, where $s^{*}$ represents the ground truth of SDF. Hand joint loss is $\mathcal{L}_{Hand}=\left\lVert\mathcal{J}_{h}-\mathcal{J}_{h}^{*}\right\rVert_{2}^{2}$, with $\mathcal{J}_{h}^{*}$ the ground truth of hand keypoints.
4.3 Force-aware Hand-pose Optimization, FO
After proper training, the VDT-Net, as well as other visual-only approaches, can give decent outputs of hand-object reconstruction. However, information near the contact areas is lacking due to self-occlusion, and the contact details should be further improved. Here, we describe an optimization method with force readings.

### v2

@@ Abstract
We present ViTaM-D, a novel visual-tactile framework for reconstructing dynamic hand-object interaction with distributed tactile sensing to enhance contact modeling. Existing methods, relying solely on visual inputs, often fail to capture occluded interactions and object deformation. To address this, we introduce DF-Field, a distributed force-aware contact representation leveraging kinetic and potential energy in hand-object interactions. ViTaM-D first reconstructs interactions using a visual network with contact constraint, then refines contact details through force-aware optimization, improving object deformation modeling. To evaluate deformable object reconstruction, we introduce the HOT dataset, featuring 600 hand-object interaction sequences in a high-precision simulation environment. Experiments on DexYCB and HOT datasets show that ViTaM-D outperforms state-of-the-art methods in reconstruction accuracy for both rigid and deformable objects. DF-Field also proves more effective in refining hand poses and enhancing contact modeling than previous refinement methods. The code, models, and datasets are available at https://sites.google.com/view/vitam-d/.
footnotetext: * indicates equal contributions. §Cewu Lu is the corresponding author.
@@ 1 Introduction
Humans manipulate objects using both visual and tactile feedback intuitively. Modeling and reconstructing this hand-object interaction can empower many downstream tasks, including VR/AR, robotic imitation learning, and human behavior understanding. Existing methods for dynamic hand-object interaction reconstruction [33, 36, 4, 29, 12] primarily rely on vision to recover the global geometry and estimate hand-object poses. However, these methods struggle to capture the occluded contact details, particularly the object deformation at the contact interface.
@@
Figure 1: (a): The relationship between tactile information and contact geometry. Grasping different bowls with the same hand poses shows that distributed tactile arrays can capture occlusion contacts and object states. (b): Different types of distributed tactile sensors proposed in previous works (1). [18], (2). [38], (3). [26].
Recent advances in tactile sensing [26, 18, 39, 23] inspire the development of visual-tactile approaches for hand-object reconstruction tasks [31, 25, 30] to leverage fine-grained perceptual signals from contact regions. (Fig. 1.a). Among various tactile sensor types, distributed tactile sensors (Fig. 1.b) [18, 38, 26], in form factors of wearable gloves or attachable sensor arrays, show great promise in human manipulation data collection comparing to other counterparts [27, 39, 16, 23]. A previous work, ViTaM [15] proposed a novel high-density tactile glove system that is capable of generating high-accuracy force data. Meanwhile, ViTaM utilized a transformer module to directly fuse tactile sensory data with visual perception for object reconstruction, which ignores the unbalanced information between forces and point cloud features. Therefore, how to effectively integrate distributed tactile sensors and state-of-the-art visual perception tools for refined hand–object reconstruction remains underexplored.
@@
Since distributed tactile sensors can conform closely to the hand while maintaining the hand’s geometry and natural dexterity during manipulation, we can use an off-the-shelf visual approach to reconstruct the hand-object state and integrate the tactile information to refine the results. In light of this, we propose a novel Distributed Force-aware contact representation, DF-Field, and a Visual-Tactile Manipulation reconstruction framework with Distributed tactile sensing, ViTaM-D. DF-Field models the contact by considering both kinetic and potential energy in hand-object interaction. This approach enables accurate modeling of object deformation. Using DF-Field, ViTaM-D initially reconstructs the hand-object interaction with visual observations and contact state constraints by the proposed visual dynamic tracking network, and then refines the contact details with DF-Field via a force-aware hand-pose optimization process. The proposed framework leverages advances in visual-only hand-object reconstruction methods [36, 4, 29, 12] while seamlessly integrating the tactile information into an existing motion capture or estimation system to enhance the fidelity of contact modeling and overall reconstruction quality.
@@
To build and train such a visual-tactile framework, large-scale datasets that capture diverse hand-object interactions with precise tactile readings are required, especially for deformable objects. However, existing datasets for hand-object interactions [3, 35, 9] only cover rigid or articulated object manipulation and generally lack accurate tactile data. Therefore, alongside using the public dataset (e.g. DexYCB [3]) to benchmark on rigid objects with common baselines [6, 4], we also introduced a new dataset, HOT dataset, to fully benchmark our method on deformable object reconstruction. This dataset is built with ZeMa [8], a high-precision physics-based simulation environment that supports penetration-free frictional contact modeling with finite element method (FEM) to model the object deformation. The HOT dataset contains 600 sequences of hand-object interaction, with 30 deformable objects from 5 different categories and 8 camera views for each sequence.
@@
To evaluate the method, we compare our proposed ViTaM-D with previous state-of-the-art methods gSDF [6], HOTrack [4] on DexYCB, and ViTaM[15] on our proposed HOT dataset. Results show that our method significantly improves the reconstruction performance compared to previous works. We successfully demonstrated the superiority of our method in tracking both rigid and deformable objects. Besides, the effectiveness of the force-aware optimization has been validated in refining hand poses to eliminate penetrations and suboptimal contact states.
@@
(1) A visual-tactile learning framework, ViTaM-D. It contains a visual dynamic tracking network for reconstructing hand-object interactions, and a force-aware optimization process, which integrates the tactile information into reconstruction refinement based on a novel distributed force-aware contact representation, DF-Field.
@@ 2 Related Work
These works focus more on visual-only inputs. However, due to occlusion between the hand and object during the interaction, visual perception usually lacks information near the contact areas, and such information cannot necessarily be mitigated by cross-frame feature fusion. Therefore, tactile perception comes to supplement the near-contact information. Zhang et al. [41] utilized a tactile glove [26] to track object movements. However, it focuses more on the dynamic object trajectory rather than contact geometry. Works such as [25, 30] employed optical tactile sensors, trained the model by synthetic data, and applied to rigid object geometry reconstruction. Later, VTacO [31] extended this line of research by using optical tactile sensors to capture object geometry, including deformation, providing a more comprehensive representation of hand-object interactions. Jiang et al. [15] proposed a visual-tactile recording and tracking system ViTaM, which contains a wearable tactile glove along with a visual-tactile reconstruction algorithm. However, it simply used neural networks to fuse tactile and visual data and did not consider the information mismatch between them. Unlike these works, the proposed ViTaM-D inherits the merits from both worlds: it leverages the advanced techniques of visual reconstruction and incorporates the distributed tactile readings to enhance the local contact details.
@@ 3 Force-aware Contact Representation
During manipulation, forces are reciprocal, so the force exerted by the hand affects the state of both the hand and the object mutually. To fully capture dynamic contact behaviors, the representation must encode both contact locations and forces. We introduce a distributed force-aware contact representation, DF-Field (Sec. 3.1), and apply it using distributed tactile sensors (Sec. 3.2).
@@ 3.1 DF-Field Representation
Without loss of generality, we take the object-centric perspective to describe a hand-object interaction process, where the object is assumed to be fixed in the origin point, and the hand moves around the object. In this way, by ignoring the gravitational potential energy, the rest of the contact dynamics is the kinetic energy and deformation potential energy of the hand and object, which are described by a Relative Potential Energy. Besides, to prevent penetration between the hand and object contact surface, we assume a virtual Barrier Energy. With an established point pair $i,j$ of object vertex $\mathcal{V}_{i}^{o}$ and hand vertex $\mathcal{V}_{j}^{h}$, and the Euclidean distance of the point pair $l_{ij}=\left\lVert\mathcal{V}_{i}^{o}-\mathcal{V}_{j}^{h}\right\rVert_{2}$, we define the two energy terms as follows.
@@
Adopted from [17], the barrier energy aims to push away the point pair when $\kappa$ is low, thus avoiding penetration issues between the hand and object. The function is defined in this way not only to ensure repulsion becomes smaller when the distance is larger but also to remain smooth for optimization.
@@
To note, though $\kappa$ is strongly correlated to the tactile readings, DF-Field can also be tactile-independent, as long as we empirically set the exerted force instead of physics-based. In this way, DF-Field can work with visual-only methods. The details are discussed in Sec. 6.5.
Figure 2: (a): The 22 regions and a typical distributed tactile sensor layout [26]. (b): An example of data collection in the simulation environment, generating depth images and tactile sensor readings.
@@ 3.2 DF-Field with Distributed Tactile Sensors
For optimization, we connect only the hand keypoints to object vertices, reducing computational load while enhancing regional interaction information. The force in each region is calculated by averaging the tactile readings $\mathcal{M}^{j}$. By the definition of the energy, we can obtain that $\kappa$ is the difference of the exerted force over the distance $l_{ij}$. In practice, we approximate it by dividing the force by the distance:
@@
We discuss the difference between using keypoints or all-hand vertices for optimization in the supplementary file.
@@ 4 ViTaM-D Method
Figure 3: ViTaM-D pipeline. (a): The overview of our pipeline, including the Flow Prediction module for visual feature extraction and flow estimation, Hand and Object Decoders, the Contact Constraint, and Force-aware hand-pose Optimization. (b): The Object Decoder to reconstruct the object mesh. (c): The Hand Decoder for estimating hand parameters and reconstructing based on the MANO model.
@@ 4.1 Overview
We address the problem of 4D tracking and dynamic reconstruction of hand-object interactions from a visual-tactile perspective. Based on DF-Field, we propose a novel method named ViTaM-D, which consists of two stages: visual dynamic tracking and force-aware optimization. We assume that the visual input consists of a live stream of 3D point clouds, denoted as $\{\mathcal{P}_{t}\in\mathbb{R}^{N\times 3}\}_{t=1}^{n}$, representing hand-object interactions derived from single-view depth images, where $N$ is the number of input point, $n$ is the number of frames. The tactile input $\{\mathcal{M}_{t}\}_{t=1}^{n}$ contains distributed tactile sensor readings for each frame.
In the first stage, ViTaM-D uses visual input to establish flow features and dynamically reconstruct the hand and object. Contact information is used to constrain the reconstruction to capture global hand-object geometry. In the second stage, tactile information is incorporated to improve contact accuracy and reduce penetration issues based on our proposed DF-Field. The overall pipeline is given in Fig. 3.
4.2 Visual dynamic tracking of hand and object
In the first stage, to introduce correspondence feature from the previous frame, the visual dynamic tracking net predicts the point cloud flow from the last to the current frame using a flow prediction module, and extracts a fused visual feature containing both correspondence feature and static information. This feature is forwarded into an object decoder and a hand decoder to reconstruct the object and hand geometry, respectively. To improve global geometry accuracy, we introduce a novel contact constraint that reasons about object geometry based on hand reconstruction. Objects are modeled using signed distance fields (SDF), while the hand is represented using the MANO model [24].
@@ 4.2 Visual dynamic tracking of hand and object
The flow prediction module consists of several MLPs and convolutional networks. After obtaining the point cloud flow $f_{t-1\rightarrow t}$, we use another PointNet++ $I^{\prime}(\cdot)$ to extract the visual correspondence feature $F_{t}^{f}$ from $\mathcal{P}_{t-1}$ with $f_{t-1\rightarrow t}$ added. $F_{t}^{f}$ and $F_{t}$ are then forwarded to a transformer fusion module to obtain the final visual feature $F_{v}$, corresponding to the current frame’s point cloud $\mathcal{P}_{t}$. Specifically, the transformer fusion module first uses self-attention to encode the point cloud features. After applying positional embedding between the feature and its point cloud, a cross-attention module fuses the two features and outputs $F_{v}$. This fusion strategy considers both 3D static information of the current frame and the corresponding feature extracted from the flow. We will discuss the detailed architecture of the flow prediction module, the importance of it, and flow prediction results in the supplementary material.
@@
Object Decoder. The object decoder processes the final visual feature $F_{v}$ in two steps: feature scattering & sampling and SDF decoding, to predict SDF values and reconstruct the object mesh using the Marching Cubes algorithm [19]. Following ConvOccNet [21], we scatter features into a volume of resolution $D$, process them with a 3D-UNet, and predict signed distances using a 5-layer MLP.
Hand Decoder. For hand reconstruction, we use the parametric MANO model [24], with $\beta\in\mathbb{R}^{10}$ for shape and $\theta\in\mathbb{R}^{51}$ for pose. Based on $F_{v}$ and the previous frame’s point cloud $\mathcal{P}_{t-1}$, we predict hand joint positions $\mathcal{J}_{h}\in\mathbb{R}^{21\times 3}$ using a voting mechanism inspired by PVN3D [14]. This involves predicting a translation offset $\mathcal{O}_{t}\in\mathbb{R}^{21\times 3}$ and a probability matrix $T_{t-1}\in\mathbb{R}^{N\times 21}$ to compute:
$\mathcal{J}_{h}=\mathcal{O}_{t}+\mathcal{P}_{t-1}\times T_{t-1}.$ (6)
Joint poses $\theta$ are then estimated via inverse kinematics [43], and the hand mesh is reconstructed using the MANO layer.
Contact Constraint. In order to obtain a relatively reasonable contact state of the reconstructed hand and object, we use a contact constraint to better optimize the prediction of the SDF of every sample position. Assuming we have the contact information $c_{x}\in\{0,1\}$ of a sample position $x\in\mathcal{X}$, then if $c_{x}=1$, it indicates that at this position, the object is in contact with the hand, meaning its signed distance value $s_{x}$ should be 00, and otherwise not equal to 00. Therefore, the contact constraint should be as follows:
$\mathcal{L}_{C}=\sum_{x\in\mathcal{X}}s_{x}\cdot\mathds{1}_{c_{x}=1}.$ (7)
If the tactile readings $\mathcal{M}$ can be obtained by haptic sensors, we can locate regional tactile arrays $\mathcal{M}^{j}$ through the hand reconstruction result, and the contact state $c_{x}$ of sample points whose distances are less than $l_{c}$ from those active tactile sensors will be set to 1, otherwise 0. In the absence of access to tactile signals, we can also use a simple network to predict contact states, which consists of a PointNet-based classifier with the input of sample positions and the final visual feature $F_{v}$. We will discuss the different effects of tactile readings and network prediction for acquiring interaction states in Sec. 6.5.
Loss Function. The network is trained end-to-end with four loss terms:
$\mathcal{L}=\lambda_{f}\mathcal{L}_{flow}+\lambda_{S}\mathcal{L}_{SDF}+\lambda_{H}\mathcal{L}_{Hand}+\lambda_{C}\mathcal{L}_{C}.$ (8)
@@
Flow loss $\mathcal{L}_{flow}$ uses Chamfer distance between forward- and backward-shifted point clouds. SDF loss $\mathcal{L}_{SDF}=|s-s^{*}|$ compares predicted and ground truth SDF values. Hand joint loss $\mathcal{L}_{Hand}=\left\lVert\mathcal{J}_{h}-\mathcal{J}_{h}^{*}\right\rVert_{2}^{2}$ penalizes deviations from ground truth keypoints. The contact constraint $\mathcal{L}_{C}$ ensures accurate SDF predictions based on contact states.
4.3 Force-aware Hand-pose Optimization
The visual dynamic tracking net can give decent outputs of hand-object reconstruction. However, even though the contact constraint module has taken contact states into account for object reconstruction, the hand pose still needs refinement, for further improving the contact details. Here, we introduce the force-aware optimization for hand pose.